if x+(1/x)=2 then x=?
plz urgent

Answer by Theo(13342)

x + 1/x = 2
multiply both sides of this equation by x to get:
x^2 + 1 = 2x
subtract 2x from both sides of this equation to get:
x^2 - 2x + 1 = 0
this factors (x-1)^2 = 0
take the square root of both sides of this equation to get:
x-1 = 0
add 1 to both sides of this equation to get:
x = 1
that should be your answer.
let's see if it's right.
substitute 1 for x in your original equation to get:
1 + 1/1 = 2
simplify this to get:
2 = 2
x = 1 is a solution to your equation.